# On the cross-correlation properties of large-size families of Costas arrays

### v1

@@ Abstract
Costas arrays have been an interesting combinatorial object for decades because of their optimal aperiodic auto-correlation properties. Meanwhile, it is interesting to find families of Costas arrays or extended arrays with small maximal cross-correlation values, since for applications in multi-user systems, the cross-interferences between different signals should also be small. The objective of this paper is to study several large-size families of Costas arrays or extended arrays, and their values of maximal cross-correlation are partially bounded for some cases of horizontal shifts $u$ and vertical shifts $v$. Given a prime $p\geq 5$, in particular, a large-size family of Costas arrays over $\{1,\ldots,p-1\}$ is investigated, including both the exponential Welch Costas arrays and logarithmic Welch Costas arrays. An upper bound on the maximal cross-correlation of this family for arbitrary $u$ and $v$ is given. We also show that the maximal cross-correlation of the family of power permutations over $\{1,\ldots,p-1\}$ for $u=0$ and $v\neq 0$ is bounded by $\frac{1}{2}+\sqrt{p-1}$. Furthermore, we give the first nontrivial upper bound of $(p-1)/t$ on the maximal cross-correlation of the larger family including both exponential Welch Costas arrays and power permutations over $\{1,\ldots,p-1\}$ for arbitrary $u$ and $v=0$, where $t$ is the smallest prime divisor of $(p-1)/2$.
@@ I Introduction
Costas arrays, first introduced by Costas in 1965 [1], have attracted much attention for decades due to their optimal aperiodic auto-correlation properties. A Costas array of order $n$ is a permutation matrix of dots/1’s and blanks/0’s such that there exists exactly one dot in each row and each column, respectively, and its aperiodic auto-correlation values are no greater than $1$ except the trivial value $n$. The aperiodic auto-correlation value, at the horizontal shift $u$ and the vertical shift $v$, is the number of coincidences between dots in the permutation matrix and dots in its corresponding shifted version. Costas arrays play a crucial role in the detection of both the position and the velocity of target objects in radar and sonar systems, as they possess such an optimal “thumbtack” ambiguity function of auto-correlation values.
@@
While the sidelobe auto-correlation values of Costas arrays are always no greater than $1$, Costas arrays in general perform poorly in terms of their cross-correlation properties. In 1985, Freedman and Levanon [2] proved that every pair of different Costas arrays of order $n\geq 4$ has a vector in common, which implies that the maximal cross-correlation of Costas arrays of order $n$ is at least $2$. Despite this fact, it is still interesting to find families of Costas arrays with small maximal cross-correlation values, since in particular for applications in multi-user systems, the cross-interferences between different signals should be restricted to a level much smaller than the order of the arrays [3]. Note that it remains a longstanding open problem whether Costas arrays exist for an arbitrary order $n$ [4] and most of Costas arrays are sporadic ones which cannot be constructed systematically [5],[6]. The two widely investigated families of Costas arrays are the family of Welch Costas arrays and the family of Lemple-Golomb Costas arrays [7],[8],[9].
@@
Recently, there have been some new results on the two algebraically constructed families of Costas arrays. For the family of exponential Welch Costas arrays, denoted by $\mathcal{W}_{p}$, Drakakis, Gow, Rickard, Sheekey and Taylor [7] showed that its maximal cross-correlation, for arbitrary $u$ and $v=0$, equals $\frac{p-1}{t}$, where $p$ is a prime and $t$ is the smallest prime divisor of $\frac{p-1}{2}$. For the case that $u$ is arbitrary and $v\neq 0$, Gómez-Pérez and Winterhof [8] further derived the upper bound of $1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor$, using the tools of character theory and the Weil bound. These results are summarized in Table I. The recent progress on the maximal cross-correlation of the family of Lemple-Golomb Costas arrays can be found for example in [7],[8],[9].
@@
While studying the differential property of power functions, Ardalani [10] observed an interesting relation between the maximal cross-correlation of the family of exponential Welch Costas arrays and the differential property of the family of power permutations, denoted by $\mathcal{P}_{p}$. As Ardalani [10] thereby defined the cross-correlation of the family of power permutations, the maximal cross-correlation of the family $\mathcal{P}_{p}$ was partially bounded. More precisely, Ardalani [10] proved that the maximal cross-correlation of power permutations equals $\frac{p-1}{t}$ for the special case $u=v=0$, where $t$ is the smallest prime divisor of $\frac{p-1}{2}$, and is bounded by $\left\lceil\frac{p-2}{p-1}(1+\sqrt{p})\right\rceil$ for $v=0$ and $u\neq 0$. The remaining two cases $u=0,v\neq 0$ and $u\neq 0,v\neq 0$ are left as open problems. Ardalani [10] further extended the original family of exponential Welch Costas arrays $\mathcal{W}_{p}$ by including the power permutations, denoted by $\mathcal{PW}_{p}$, and examined its cross-correlation properties. However, except some numerical results, generic theoretical bounds on the maximal cross-correlation of this extended family are missing.
@@
The objective of this paper is to establish certain theoretical bounds on the maximal cross-correlation of several large-size families of Costas arrays or extended arrays, and the main contributions are summarized as follows. First, a large-size family of Costas arrays, denoted by $\mathcal{W}_{p}^{el}$, is introduced by combining together the family of exponential Welch Costas arrays and that of logarithmic Welch Costas arrays. An upper bound on the maximal cross-correlation of this large family for arbitrary $u$ and $v$ is derived, using the idea that transmits the transcendental equations into polynomial functions [11] (see Table I). We then revisit the family $\mathcal{PW}_{p}$, and give the first nontrivial upper bound of its maximal cross-correlation for arbitrary $u$ and $v=0$. The main tool is to transform the problem of calculating the cross-correlation values into that of counting the number of fixed points of Welch Costas arrays. Meanwhile, we are able to derive an upper bound on the maximal cross-correlation of the family $\mathcal{P}_{p}$ for $u=0$ and $v\neq 0$. See Table II for relevant bounds on certain cases of these large-size families of extended arrays.
@@
$\mathcal{W}_{p}$ $\phi(p-1)$ $\frac{p-1}{t}$ arbitrary u, $v=0$ [7]
$1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor$ arbitrary u, $v\neq 0$ [8]
$\mathcal{W}_{p}^{el}$ $2\phi(p-1)$ $\max\left\{4p\,\log_{p}\alpha+1,\frac{p-1}{t},1+\left\lfloor(1-\frac{2}{p-1})p^{1/2}\right\rfloor\right\}$ arbitrary $u$ and $v$ This paper
@@
$\mathcal{P}_{p}$ $\phi(p-1)-1$ $\frac{p-1}{t}$ $u=0,v=0$ [10]
$\frac{p-2}{p-1}(1+\sqrt{p})$ $u\neq 0,v=0$ [10]
@@
$\mathcal{PW}_{p}$ $12\phi(p-1)-$ 1 $\frac{p-1}{t}$ arbitrary $u$, $v=0$ This paper
@@ II Preliminaries
Let $n$ be a positive integer and $[n]$ denote the set $\{1,2,\ldots,n\}$. Let $u,v$ be two nonnegative integers, which typically denote the horizontal and the vertical shifts of a permutation matrix, respectively. In the following, we give the formal definitions of the aperiodic auto-correlation and cross-correlation of Costas arrays.
@@
Let $f=[f(i)]=[f(1),f(2),\ldots,f(n)]:[n]\rightarrow[n]$ be a permutation of order $n$, and let $A=(a_{ij})_{n\times n}$ be the corresponding permutation matrix of $f$: $a_{ij}=1$ if and only if $f(j)=i$. Then $A$ is a Costas array if the aperiodic auto-correlation of A satisfies $\Psi_{A,A}(u,v)\leq 1$ for all $(u,v)\neq(0,0)$, where $\Psi_{A,A}(u,v)=|\{(f(i)+v,i+u):i\in[n]\}\cap\{(f(i),i):i\in[n]\}|$.
@@
For $-5\leq u,v\leq 5$, we calculate the auto-correlation value $\Psi_{A,A}(u,v)$ of $A$ at the shift $(u,v)$ relative to itself. The results are recorded in the matrix $\mathcal{C}_{A}$ as follows, with the $(0,0)$ shift positioned centrally.
@@
The definition above of Costas arrays is via aperiodic auto-correlation and the equivalent two definitions of Costas arrays via distance vectors and difference triangles can be found for example in [4, 12].
Definition 2.
Below we recall the classical algebraic construction of Costas arrays by Welch [12], and the other classical Lempel-Golomb construction can be found in [12]. Denote by $\mathbb{F}_{p}$ the finite field of order $p$, where $p>2$ is a prime. Let $\alpha$ be a primitive element of $\mathbb{F}_{p}$ and $c$ be an integer with $20\leq c\leq p-$ 2. A prime $p$ is a safe prime if and only if $p=2r+1$ and $r$ is also a prime (also known as Sophie Germain prime).
@@
[Exponential Welch Construction] The $(p-1)\times(p-1)$ permutation matrix, denoted by $W_{1}^{\exp}(p,\alpha,c)=(a_{ij})_{(p-1)\times(p-1)}$, where $a_{ij}=1$ if and only if $i\equiv\alpha^{j-1+c}(\bmod\,p)$, for $11\leq i,j\leq p-$ 1, is a Costas array.
@@
[Logarithmic Welch Construction] The $(p-1)\times(p-1)$ permutation matrix, denoted by $W_{1}^{\log}(p,\alpha,c)=(a_{ij})_{(p-1)\times(p-1)}$, where $a_{ij}=1$ if and only if $i\equiv 1+c+\log_{\alpha}j\bmod(p-1)$, for $11\leq i,j\leq p-$ 1, is a Costas array.
In a Costas array, the points on the main diagonal are called fixed points. The number of fixed points plays an important role in understanding the structure of the algebraically constructed Costas arrays [13, 14, 15]. In fact, the fixed points in a Costas array form a Golomb ruler (or called Sidon set, $B_{2}$-sequence). Formally, a Golomb ruler is a set of distinct integers such that all differences between any two integers are different. Therefore, the number of the fixed points in Costas arrays of order $n$ is naturally bounded by the size of the Golomb ruler in the set $[n]$. Cilleruelo [16] gave the following upper bound on the size of Golomb rulers while studying the size of Sidon sets.
@@
[16, Corollary 2] Let $F(n)$ denote the maximal cardinality of a Sidon set in $[n]$, then
@@
Note that the cross-correlation between a Costas array constructed by the exponential Welch construction (Construction 1) and its cyclically shifted version may contain values between $1$ and $p-2$, where the cyclically shifted version means a Costas array by the exponential Welch construction using the same primitive element but a different constant $c$. The family of exponential Welch Costas arrays, denoted by $\mathcal{W}_{p}$, does not contain cyclically shifted Costas arrays. In other words, the constant $c$ is fixed and set to 00 for convenience to generate a family of Costas arrays by Construction 1. In a similar way, the family of logarithmic Welch Costas arrays, denoted by $\mathcal{W}_{p}^{l}$, is also defined in the following.
@@
For prime $p\geq 5$, the family of exponential Welch Costas arrays of order $(p-1)$ that does not contain any cyclic shift of a given exponential Welch array $W_{1}^{\exp}(p,\alpha,0)$, is defined as
$\mathcal{W}_{p}=\left\{\left[\alpha^{i-1}\bmod p\right]:1\leq i\leq p-1\text{ and }\alpha\text{ is a primitive element of }\mathbb{F}_{p}\right\}.$
@@
For prime $p\geq 5$, the family of logarithmic Welch Costas arrays of order $(p-1)$ that does not contain any cyclic shift of a given logarithmic Welch array $W_{1}^{\log}(p,\alpha,0)$, is defined as
$\mathcal{W}^{l}_{p}=\left\{\left[(1+\log_{\alpha}i)\bmod(p-1)\right]:1\leq i\leq p-1\text{ and }\alpha\text{ is a primitive element of }\mathbb{F}_{p}\right\}.$
@@
[17, Theorem 5.41] Let $\mathbb{F}_{p}$ be a finite field with $p$ elements and let $\psi$ be a multiplicative character of $\mathbb{F}_{p}$ of order $m>1$. Suppose that $f\in\mathbb{F}_{p}[x]$ has precisely $d$ distinct roots in its splitting field over $\mathbb{F}_{p}$, and suppose that $f$ is not an m-th power of a polynomial. Then for every $a\in\mathbb{F}_{p}$, we have
@@
The Weil bound above becomes trivial in some cases, for example, when the number of distinct roots $d$ is large. Kelley and Owen [18] gave the following upper bound on the number of roots of trinomials over finite fields, which will be used in deriving bounds on the maximal cross-correlation of Costas arrays afterwards.
@@
[…]
The family of power permutations $\mathcal{P}_{p}$, studied by Ardalani [10], is defined as follows.
@@
For prime $p\geq 5$, the family $\mathcal{P}_{p}$ of power permutations of order $(p-1)$, is defined as
$\mathcal{P}_{p}=\left\{\left[x^{d}\bmod p\right]:1\leq x\leq p-1\text{ and }\operatorname{gcd}(d,p-1)=1\right\}.$

### v2

@@ Abstract
Costas arrays have been an interesting combinatorial object for decades because of their optimal aperiodic auto-correlation properties. Meanwhile, it is interesting to find families of Costas arrays or extended arrays with small maximal cross-correlation values, since for applications in multi-user systems, the cross-interferences between different signals should also be small. The objective of this paper is to study several large-size families of Costas arrays or extended arrays, and their values of maximal cross-correlation are partially bounded for some cases of horizontal shifts $u$ and vertical shifts $v$. Given a prime $p\geq 5$, a large-size family of Costas arrays over $\{1,\ldots,p-1\}$ is investigated, including both the exponential and logarithmic Welch Costas arrays. An upper bound on the maximal cross-correlation of this family for arbitrary $u$ and $v$ is given. We also show that the maximal cross-correlation of the family of power permutations over $\{1,\ldots,p-1\}$ for $u=0$ and $v\neq 0$ is bounded by $\frac{1}{2}+\sqrt{p-1}$. Furthermore, we give the first nontrivial upper bound on the maximal cross-correlation of the larger family including both exponential Welch Costas arrays and power permutations over $\{1,\ldots,p-1\}$ for arbitrary $u$ and $v=0$ that it equals (p1)/t(p-1)/t where $t$ is the smallest prime divisor of (p1)/2(p-1)/2 if p is not a safe prime and is at most $(p-1)^{\frac{1}{2}}+(p-1)^{\frac{1}{4}}+\frac{1}{2}$ otherwise.
@@ I Introduction
Costas arrays, first introduced by Costas in 1965 [Costas1965Sonar], have attracted much attention for decades due to their optimal aperiodic auto-correlation properties. A Costas array of order $n$ is a permutation matrix of dots/1’s and blanks/0’s such that there exists exactly one dot in each row and each column, respectively, and its aperiodic auto-correlation values, except for the trivial case, are less than or equal to $1$. The aperiodic auto-correlation value, at the horizontal shift $u$ and the vertical shift $v$, is the number of coincidences between dots in the permutation matrix and dots in its corresponding shifted version. Costas arrays play a crucial role in the detection of both the position and the velocity of target objects in radar and sonar systems, as they possess such an optimal “thumbtack” ambiguity function of auto-correlation values [GT1982Two],[Costas1984study].
@@
While the auto-correlation sidelobe values of Costas arrays are always no greater than $1$, Costas arrays in general perform poorly in terms of their cross-correlation properties. In 1985, Freedman and Levanon [FL1985Two] proved that the maximal cross-correlation of every pair of different Costas arrays of order $n\geq 4$ is at least $2$. Despite this fact, it is still interesting to find families of Costas arrays with small maximal cross-correlation values, since in particular for applications in multi-user systems, the cross-interferences between different signals should be restricted to a level much smaller than the order of the arrays [DT1991Cross-correlation]. Note that it remains a longstanding open problem whether Costas arrays exist for an arbitrary order $n$ [GT1984Constructions],[WCS2023] and most of Costas arrays are sporadic ones which cannot be constructed systematically [AP2022Transformation, TR2011Matlab]. The two widely investigated families of Costas arrays are the family of Welch Costas arrays and the family of Lempel-Golomb Costas arrays [DG2011Maximal, GW2020Note, LW2024cross-correlation].
@@
Recently, there have been some new results on the two algebraically constructed families of Costas arrays. For the family of exponential Welch Costas arrays, denoted by $\mathcal{W}_{p}$, Drakakis, Gow, Rickard, Sheekey and Taylor [DG2011Maximal] showed that its maximal cross-correlation, for arbitrary $u$ and $v=0$, equals $\frac{p-1}{t}$, where $p$ is a prime and $t$ is the smallest prime divisor of $\frac{p-1}{2}$. For the case that $u$ is arbitrary and $v\neq 0$, Gómez-Pérez and Winterhof [GW2020Note] further derived the upper bound of $1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor$, using the tools of character theory and the Weil bound. These results are summarized in Table I. The recent progress on the maximal cross-correlation of the family of Lempel-Golomb Costas arrays can be found for example in [DG2011Maximal, GW2020Note, LW2024cross-correlation].
@@
While studying the differential property of power functions, Ardalani [Ardalani2023Contributions] observed an interesting relation between the maximal cross-correlation of the family of exponential Welch Costas arrays and the differential property of the family of power permutations, denoted by $\mathcal{P}_{p}$. As Ardalani thereby defined the cross-correlation of the family of power permutations, the maximal cross-correlation of the family $\mathcal{P}_{p}$ was partially bounded. More precisely, Ardalani proved that the maximal cross-correlation of power permutations equals $\frac{p-1}{t}$ for the special case $u=v=0$, where $t$ is the smallest prime divisor of $\frac{p-1}{2}$, and is bounded by $\left\lceil\frac{p-2}{p-1}(1+\sqrt{p})\right\rceil$ for $v=0$ and $u\neq 0$. The remaining two cases $u=0,v\neq 0$ and $u\neq 0,v\neq 0$ are left as open problems. Ardalani further extended the original family of exponential Welch Costas arrays $\mathcal{W}_{p}$ by including the power permutations, denoted by $\mathcal{PW}_{p}$, and examined its cross-correlation properties. However, except some numerical results, generic theoretical bounds on the maximal cross-correlation of this extended family are missing.
@@
The objective of this paper is to establish certain theoretical bounds on the maximal cross-correlation of several large-size families of Costas arrays or extended arrays, and the main contributions are summarized as follows. First, a large-size family of Costas arrays, denoted by $\mathcal{W}_{p}^{el}$, is introduced by combining together the family of exponential Welch Costas arrays and that of logarithmic Welch Costas arrays. An upper bound on the maximal cross-correlation of this large family for arbitrary $u$ and $v$ is derived, using the idea that transforms the transcendental equations into polynomial functions [GS2010Cycles] (see Table I). We then revisit the family $\mathcal{PW}_{p}$, and give the first nontrivial upper bound of its maximal cross-correlation for arbitrary $u$ and $v=0$. The main tool is to transform the problem of calculating the cross-correlation values into that of counting the number of fixed points of Welch Costas arrays. Meanwhile, we are able to derive an upper bound on the maximal cross-correlation of the family $\mathcal{P}_{p}$ for $u=0$ and $v\neq 0$. See Table II for relevant bounds on certain cases of these large-size families of extended arrays.
@@
$\mathcal{W}_{p}$ $\phi(p-1)$ $\frac{p-1}{t}$ arbitrary u, $v=0$ [DG2011Maximal]
$1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{\frac{1}{2}}\right\rfloor$ arbitrary u, $v\neq 0$ [GW2020Note]
$\mathcal{W}_{p}^{el}$ $2\phi(p-1)$ $\max\left\{4p\log_{p}\alpha+1,1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{\frac{1}{2}}\right\rfloor\right\}$, if $p$ is a safe prime arbitrary $u$, $v$ This paper
$\max\left\{4p\log_{p}\alpha+1,\frac{p-1}{t}\right\}$, if $p$ is not a safe prime
@@
$\mathcal{P}_{p}$ $\phi(p-1)-1$ $\frac{p-1}{t}$ $u=0,v=0$ [Ardalani2023Contributions]
$\frac{p-2}{p-1}(1+\sqrt{p})$ $u\neq 0,v=0$ [Ardalani2023Contributions]
@@
$\mathcal{PW}_{p}$ $2\phi(p-1)-1$ $(p-1)^{\frac{1}{2}}+(p-1)^{\frac{1}{4}}+\frac{1}{2}$, if $p$ is a safe prime arbitrary $u$, $v=0$ This paper
$\frac{p-1}{t}$, if $p$ is not a safe prime
@@ II Preliminaries
Let $n$ be a positive integer and $[n]$ denote the set $\{1,2,\ldots,n\}$. Let $u,v$ be two integers, which typically denote the horizontal and the vertical shifts of a permutation matrix, respectively. A bijection $f$ from $[n]$ to $[n]$ is called a permutation of order $n$, denoted as $f=[f(i)]=[f(1),f(2),\ldots,f(n)]$. The corresponding permutation matrix is denoted by $A_{f}=(a_{ij})_{n\times n}$ where $a_{ij}=1$ if and only if $f(j)=i$ and $a_{ij}=0$ otherwise. When no ambiguity arises, we omit subscript $f$. This permutation matrix can be regarded as an equivalent representation of the permutation $f$.
@@
In particular, the cross-correlation is the aperiodic auto-correlation when $f=g$. In the following, we give the formal definitions of Costas arrays through the property of the aperiodic auto-correlation.
Definition 2.
Let $f=[f(i)]=[f(1),f(2),\ldots,f(n)]:[n]\rightarrow[n]$ be a permutation of order $n$, and let $A=(a_{ij})_{n\times n}$ be the corresponding permutation matrix. Then $A$ is a Costas array if the aperiodic auto-correlation of A satisfies $\Psi_{A,A}(u,v)\leq 1$ for all $(u,v)\neq(0,0)$.
@@
For $5\leq u,v\leq 5$, we calculate the auto-correlation value $\Psi_{A,A}(u,v)$ of $A$ at the shift $(u,v)$ relative to itself. The results are recorded in the matrix $\mathcal{C}_{A}$ as follows, with the $(0,0)$ shift positioned centrally.
@@
The definition above of Costas arrays is via aperiodic auto-correlation and the equivalent two definitions of Costas arrays via distance vectors and difference triangles can be found for example in [GT1984Constructions, Golomb1984Algebraic].
Below we recall the classical algebraic construction of Costas arrays by Welch [Golomb1984Algebraic], and the other classical Lempel-Golomb construction can be found in [Golomb1984Algebraic].
@@
[GT1984Constructions] [Exponential Welch Construction] Denote by $\mathbb{F}_{p}$ the finite field of order $p$, where $p>2$ is a prime. Let $\alpha$ be a primitive element of $\mathbb{F}_{p}$ and $c$ be an integer with $0\leq c\leq p-2$. The $(p-1)\times(p-1)$ permutation matrix, denoted by $W_{1}^{\exp}(p,\alpha,c)=(a_{ij})_{(p-1)\times(p-1)}$, where $a_{ij}=1$ if and only if $i\equiv\alpha^{j-1+c}(\bmod\,p)$, for $1\leq i,j\leq p-1$, is a Costas array.
@@
[Logarithmic Welch Construction] Denote by $\mathbb{F}_{p}$ the finite field of order $p$, where $p>2$ is a prime. Let $\alpha$ be a primitive element of $\mathbb{F}_{p}$ and $c$ be an integer with $0\leq c\leq p-2$. The $(p-1)\times(p-1)$ permutation matrix, denoted by $W_{1}^{\log}(p,\alpha,c)=(a_{ij})_{(p-1)\times(p-1)}$, where $a_{ij}=1$ if and only if $i\equiv 1+c+\log_{\alpha}j\bmod(p-1)$, for $1\leq i,j\leq p-1$, is a Costas array.
In a Costas array, the points on the main diagonal are called fixed points. The number of fixed points plays an important role in understanding the structure of the algebraically constructed Costas arrays [Drakakis2011Open, JW2014Deficiency, JY2018cubes]. In fact, the fixed points in a Costas array form a Golomb ruler (or called Sidon set, $B_{2}$-sequence). Formally, a Golomb ruler is a set of distinct integers such that all differences between any two integers are different. Therefore, the number of the fixed points in Costas arrays of order $n$ is naturally bounded by the size of the Golomb ruler in the set $[n]$. Cilleruelo [Cilleruelo2010sidon] gave the following upper bound on the size of Golomb rulers while studying the size of Sidon sets.
@@
[Cilleruelo2010sidon, Corollary 2] Let $F(n)$ denote the maximal cardinality of a Sidon set in $[n]$, then
@@
Note that the cross-correlation between a Costas array constructed by the exponential Welch construction (Construction 1) and its cyclically shifted version may contain values between $1$ and p2p-2, where the cyclically shifted version means a Costas array by the exponential Welch construction using the same primitive element but a different constant $c$. The family of exponential Welch Costas arrays, denoted by $\mathcal{W}_{p}$, does not contain cyclically shifted Costas arrays. In other words, the constant $c$ is fixed and set to 0 for convenience to generate a family of Costas arrays by Construction 1. In a similar way, the family of logarithmic Welch Costas arrays, denoted by $\mathcal{W}_{p}^{l}$, is also defined in the following.
@@
For prime $p\geq 5$, the family of exponential Welch Costas arrays of order (p1)(p-1) that does not contain any cyclic shift of a given exponential Welch array $W_{1}^{\exp}(p,\alpha,0)$, is defined as
[…]
@@
For prime $p\geq 5$, the family of logarithmic Welch Costas arrays of order (p1)(p-1) that does not contain any cyclic shift of a given logarithmic Welch array $W_{1}^{\log}(p,\alpha,0)$, is defined as
[…]
@@
[LN1997finitefield, Theorem 5.41] Let $\mathbb{F}_{p}$ be a finite field with $p$ elements and let $\psi$ be a multiplicative character of $\mathbb{F}_{p}$ of order $m>1$. Suppose that $f\in\mathbb{F}_{p}[x]$ has precisely $d$ distinct roots in its splitting field over $\mathbb{F}_{p}$, and suppose that $f$ is not an m-th power of a polynomial. Then for every $a\in\mathbb{F}_{p}$, we have
@@
The Weil bound above becomes trivial in some cases, for example, when the number of distinct roots $d$ is large. Kelley and Owen [KO2017trinomials] gave the following upper bound on the number of roots of trinomials over finite fields, which will be used in deriving bounds on the maximal cross-correlation of Costas arrays afterwards.
@@
with $a$ and $b$ both nonzero, the number of roots $R(f)\leq\delta\left\lfloor\frac{1}{2}+\sqrt{\frac{p-1}{\delta}}\right\rfloor$, where $\delta=\operatorname{gcd}(n,s,p-1)$, and $\lfloor\cdot\rfloor$ denotes the floor function.
The family of power permutations $\mathcal{P}_{p}$, studied by Ardalani [Ardalani2023Contributions], is defined as follows.
@@
For prime $p\geq 5$, the family $\mathcal{P}_{p}$ of power permutations of order (p1)(p-1), is defined as
$\mathcal{P}_{p}=\left\{\left[i^{d}\bmod p\right]:1\leq i\leq p-1\text{ where }\operatorname{gcd}(d,p-1)=1\text{ and }1<d\leq p-2\right\}.$